A company has a warehouse, it currently deals with 10 items A-J Item Quantity per Receipt 80 pallets 50 pallets Trips to ReceShipping Receiving Location 1: Location 2: Location 3 Location 4: Location 5:Location 6: Location 7: Location 8: Location : Location 10:
Answer #1

a.

Qty/receipt units are pallets for items A to E and cartons for items F to J

Avg customer order size units are pallets for items A to E and cartons for items F to J

Now trips to ship = Qty/receipt divided by Avg customer order size

For item A, this is = 80/1 = 80

Further, receiving/shipping ration = trips to receive/trips to ship = 40/80 = 0.5 for item A

Item Qty/
receipt
Trips to receive Avg cust order size Trips to ship Receiving/Shipping ration
A 80 40 1 80 0.50
B 50 100 0.4 125          0.80
C 60 200 1.5 40          5.00
D 70 60 0.7 100          0.60
E 10 5 0.4 25          0.20
F 1000 50 4 250          0.20
G 1000 200 8 125          1.60
H 2000 150 4 500          0.30
I 2000 50 40 50          1.00
J 3000 75 20 150          0.50

The table is filled in a similar way

For example for item G, trips to ship = 1000 cartons/8 = 125

Also receiving/shipping ration = 200/125 = 1.60

b. Items in the warehouse to minimize movement in the warehouse:

Shipping Receiving

To allocate items to above layout, we should look at receiving/shipping ration that will indiciate which item to be placed where

Based on above table obtained, the revised table arranged in ascending order or the ratio is shown below:

Item Receiving/Shipping ration
E          0.20
F          0.20
H          0.30
A 0.50
J          0.50
D          0.60
B          0.80
I          1.00
G          1.60
C          5.00

Looking at above table, lesser the ratio of receiving/shipping, closer the item to be placed to the shipping side of the layout i.e. on the left side

So E and F can be put in location 1 and 10 respectively

H can be placed in location 2

either A or J can be placed in location 9. However since J has higher absolute no of trips to ship i.e. 150 we will place J in location 9

So A can be placed in location 3

D can be placed in location 8

B can be placed in locaiton 4

I can be placed in location 7

G can be placed in location 5

and C in location 6